A study was done to determine the average number of homes that a homeowner owns in his or her lifetime. For the 60 homeowners surveyed, the sample average was 4.2 and the sample standard deviation was 2.1. Calculate the 95% confidence interval for the true average number of homes that a person owns in his or her lifetime.
a. (3.90, 4.50)
b. (3.66, 4.74)
c. (4.01, 4.39)
d. (3.67, 4.73)
9. A study was done to determine the average number of homes that a homeowner owns in his or her lifetime. For the 60 homeowners surveyed, the sample average was 4.2 and the sample standard deviation was 2.1. The distribution to use to calculate the 95% confidence interval is ________________.
a. Exponential
b. Student-t with df = 59
c. Student-t with df = 60
d. Binomial
10. Calculate the error bound for a survey in which the sample mean is 5, the population standard deviation is 3, the confidence level is 0.90, and the number surveyed is 12.
a. 1.56
b. 1.42
c. (3.58, 6.42)
d. (3.44, 6.56)

Answer by Boreal(15235)

interval for 95% CI is t df=59,0.975*s/sqrt(n)=2.00*2.1/sqrt(60)=0.54
4.2+/-0.54=(3.66, 4.74)
b
t df=59
This is z, so 1.645 is used
1.645*3/sqrt(12) and half-interval width is 1.42
But question asks for error bound, not error size
Answer is c (3.58, 6.42)